# Mildly Accurate Computationally Differentially Private Inner Product Protocols Imply Oblivious Transfer

### v1

@@ 2 Proof Overview
In the following, let $X,Y$ be inputs of $\mathsf{A}$ and $\mathsf{B}$ to $\Pi$, respectively, and let $V_{A}$ and $V_{B}$ be their view in the interaction according to $\Pi$. Let $R,I_{1},\dots,I_{k},\widetilde{Y}$ and $\widehat{Y}$ be the random variables taking the values of $r,i_{1},\dots,i_{k},\widetilde{y}$ and $\widehat{y}$ respectively. Finally, let $\widehat{V}_{A}=(V_{A},R,\widehat{Y}_{-R})$ and $\widehat{V}_{B}=(V_{B},R,X_{R},I_{1},\dots,I_{k},\widetilde{Y},\widehat{Y})$ be the parties’ views in 2.0.0.2, and $O_{A},O_{B}$ their outputs, respectively.
It is clear by the definition of the protocol that erasure indeed occurs with probability $1/2$. Moreover, similarly to the correctness of the key-agreement protocol, and by the accuracy of $\Pi$, when there is no erasure, $O_{A}$ and $O_{B}$ are of distance at most $\ell$ apart with probability at least $0.999$. Thus, it remains to prove that $\mathsf{B}$ cannot approximate $O_{A}$ when erasure occurs and that $\mathsf{A}$ cannot learn whether it occurred. To simplify the analysis in this proof overview, in the following, we assume that $\Pi$ is a $(\varepsilon,0)$-CDP protocol, [^5] While $(\varepsilon,0)$-CDP (that is, setting $\delta=0$) doesn’t make sense in the computational setting, for the sake of simplicity, we ignore this subtly in this informal overview. for a small constant $\varepsilon>0$.
@@
Before discussing the security proof of our protocol, we want to remark an important difference from the analysis of [HaitnerMST22]’s ${\sf KA}$ protocol. While the primitive we construct here, ${\sf OT}$, is a stronger primitive than ${\sf KA}$, some parts of the proof actually became easier in our analysis. This is due the fact that in ${\sf OT}$, we need to consider an internal attacker, while in ${\sf KA}$ the security needs to hold against external observer.
In more detail, [HaitnerMST22] needed to show that security holds against an adversary who sees $R,X_{R}$ and $Y_{-R}$. This form of leakage creates a dependency between the seed $R$ to both the leakage on $X$, $X_{R}$, and the leakage of $Y$, $Y_{-R}$. This form of dependency makes the analysis challenging. In contrast, in our setting, the attacker sees one of the vectors $X$ and $Y$ entirely. For example, $\mathsf{A}$ gets to see $X,R$ and $Y_{-R}$, and therefore there is no dependency between the leakage on $X$ and $R$. This fact helps us to bypass this dependency issue, and makes this part of our analysis easier.
2.0.0.4 Erasures do erase.
This parts follow quite easily using ideas from [McGregorMPRTV10, HaitnerMST22]. Assume towards a contradiction that there exists an efficient $\widehat{\mathsf{B}}$, that given $\widehat{V}_{B}$ and conditioned on $O_{B}=\bot$ (equivalently, on $b=1$) can approximate $O_{A}$ within small distance. Namely, assume that
@@
To see why, let ~=$\widetilde{{\cal H}}=\mathopen{}\mathclose{{}\left\{i\colon R_{i}=0}\right\}\setminus{\cal H}$, and let $T=\widehat{\mathsf{B}}(\widehat{V}_{B})-\langle X_{\widetilde{{\cal H}}},\widetilde{Y}_{\widetilde{{\cal H}}}\rangle$. We claim that $T$ is a good approximation of $\langle X_{{\cal H}},S\rangle$. Indeed,
$T-\langle X_{{\cal H}},S\rangle$ $=T-(\langle X_{-R},\widetilde{Y}_{-R}\rangle-\langle X_{\widetilde{{\cal H}}},\widetilde{Y}_{\widetilde{{\cal H}}}\rangle)$
$=(T+\langle X_{\widetilde{{\cal H}}},\widetilde{Y}_{\widetilde{{\cal H}}}\rangle)-\langle X_{-R},\widetilde{Y}_{-R}\rangle$
@@
Thus we only need to show an algorithm $\widetilde{\mathsf{B}}$, that given $V_{B},{\cal H},S,X_{-{\cal H}}$ as input, evaluates the value of $T$. The main issue left is that while $\widetilde{\mathsf{B}}$ only gets $V_{B}$, and to evaluate $T$ we need to execute $\widehat{B}$ that needs to get an entire view $\widehat{V}_{B}$ of $\mathsf{B}$ in 2.0.0.2. However, this can be easily simulated. Indeed, $\widetilde{\mathsf{B}}$ can simply sample $R,I_{1},\dots,I_{k}$ and $\widetilde{Y}$ condition on the values of ${\cal H},S$ (and $O_{B}=\bot$), to get the exact distribution of $\widehat{V}_{B}$.
2.0.0.5 Erasures are hidden.
This part is more involved and is our main new technical part. The goal in this part is to show that $\widehat{\mathsf{A}}$ cannot tell when an erasure happens. Equivalently, we want to show that $\widehat{\mathsf{A}}$ cannot distinguish $Y_{R}$ from the noisy version $\widetilde{Y}_{R}$ with too good probability. For this end, we show that an algorithm $\widehat{\mathsf{A}}$ that (given $V_{A}$) can distinguish $Y_{R}$ from the noisy version $\widetilde{Y}_{R}$ with a small but constant probability, can be used to break ${\sf CDP}$.
The first step in the proof is to notice that, by a rather standard hybrid argument, it is enough to show that $\widehat{\mathsf{A}}$ cannot distinguish with advantage $\Theta(1/k)$ between $Y$ to a vector $Y^{\prime}$ which is identical in all but one random coordinate. That is, let $I\leftarrow[n]$ be a random index, and let $Y^{(I)}=(Y_{1},\dots,Y_{I-1},-Y_{I},Y_{I+1},...,Y_{n})$ be the vector that is identical to $Y$ in all entries except the $I$-th one. Then it is enough to show that if
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A},R,Y_{R})=1}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A},R,Y^{(I)}_{R})=1}\right]\geq 1/k,$
@@
We highlight that the key difficulty in proving the above, lies in the fact that the distinguishing advantages is very small. Indeed, an algorithm $\widehat{A}$ that distinguishes $Y_{R}$ from $Y^{(I)}_{R}$ with a large enough constant advantage (proportional to $\varepsilon$) already contradicts the privacy requirement of $\Pi$. However, in our setup the distinguishing advantage is much smaller than $\varepsilon$, which makes the task of contradicting $\varepsilon$-CDP challenging. Moreover, we want to emphasize the following two crucial points that makes $\widehat{\mathsf{A}}$ useful (and our design for our AWEC protocol, 2.0.0.2, was guided by these two observations).
  • The importance of a random index. We note that it is crucial that the index $I$ is chosen at random and that it is unknown to $\widehat{\mathsf{A}}$. To see that, let $\mathsf{M}$ be a DP mechanism and $Y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ be an input. We claim that it can be the case that an adversary can distinguish between $Y$ and $Y^{(i)}$ given $\mathsf{M}(Y)$ with advantage $\approx\varepsilon$, for any fixed index $i$. Indeed, it is possible that $\mathsf{M}(Y)$ outputs a noisy estimation of $Y_{i}$, which is correct with probability $\approx 1/2+\varepsilon$. Such an estimation will not contradict differential privacy, but will be useful to distinguish between $Y$ and $Y^{(i)}$. We resolve this issue by adding the noise in a random position.
  • The importance of a random subset. Next, we claim that it is crucial that $\widehat{\mathsf{A}}$ only gets a random subset of $Y$, $Y_{R}$. Indeed, as in the previous example, the mechanism $\mathsf{M}(Y)$ can output a noisy estimation of the (parity of) the number of ones in $Y$. As in the previous example, such an estimation does not contradict DP, but is enough to distinguish $Y$ and $Y^{(I)}$ by considering the number of ones in the input compared to the estimation. We deal with such a global information by revealing to $\widehat{\mathsf{A}}$ only subset of $Y$. This makes the global information not useful.
Thus, in the proof, we need to leverage the fact that $\widehat{\mathsf{A}}$ distinguishes $Y_{R}$ from $Y^{(I)}_{R}$ for a random index $I$ and a random subset $R$. We use such an algorithm $\widehat{\mathsf{A}}$ to construct an algorithm $\mathsf{G}$ that given $Y_{-I}$, predicts the value of $Y_{I}$. In more detail, we prove the following lemma, which is the technical core of our proof.
@@
Let $(Z,W)$ be jointly distributed random variables, let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $I\leftarrow[n]$. Let $\mathsf{A}$ be an efficient algorithm that satisfy
@@
Back to our setting, using 2.0.0.5 (taking $\mathsf{A}=\widehat{\mathsf{A}},Z=Y$ and $W=V_{A}$), we get an algorithm $\mathsf{G}$ such that
@@
We highlight that the gap between the probability of $\mathsf{G}$ to output $Y_{I}$ to its probability to output $-Y_{I}$ is smaller than $1/k$, and that it is not enough to break DP alone. The point is that we have a bound on the probability of $\mathsf{G}$ to output $-Y_{I}$, and that in DP we measure the difference in probabilities with multiplicative distance. Thus, when
@@
which is the reason we need to take $k\ll n^{1/3}$. In the rest of this section we give an high level proof sketch of 2.0.0.5.
2.1 Proof Overview for 2.0.0.5
In this part we explain the ideas behind the proof of 2.0.0.5. We start with describing the algorithm $\mathsf{G}$.
@@
The construction of $\mathsf{G}$ will leverage the fact that $\mathsf{A}$ succeed on a random index $I$ and while only seeing a subset of $Y$. We next describe the algorithm $\mathsf{G}$. Recall that on input $I,Z_{-I},W$, the goal of $\mathsf{G}$ is to predict the value of $Z_{I}$. Toward this, for $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$, let $Z^{b}=(Z_{1},\dots,Z_{I-1},b,Z_{I+1},\dots,Z_{n})$, so that $Z\in\mathopen{}\mathclose{{}\left\{Z^{-1},Z^{1}}\right\}$. Then $\mathsf{G}$ computes three values: [^8] In the proof itself, $\mathsf{G}$ only estimates these values by sampling $R$. For simplicity, here we assume that $\mathsf{G}$ can compute them exactly.
@@
Importantly, we note that when $R_{I}=0$, $Z_{R}$ does not contain $Z_{I}$, and thus $\mathsf{G}$ can compute $\mu^{*}$ giving its input. Finally, $\mathsf{G}$ outputs $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ if the following two conditions holds:
@@
Therefore, assuming that Item 1 and Item 2 hold, this conclude the proof of 2.0.0.5. We next briefly explain why Item 1 and Item 2 hold with the claimed probability.
@@
In the following we will actually prove a stronger claim. That is, that for every fixed $Z$ and $W$, it holds that
@@
Finally, Equation 5 follows a more general well-known fact on the uniform distribution over random strings: For a random string $R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and for a random index $I\leftarrow[n]$, the statistical distance between $R|_{R_{I}=0}$ and $R|_{R_{I}=1}$ is at most $1/\sqrt{n}$. This fact implies that the function $F$ cannot distinguish too well between samples from $R|_{R_{I}=0}$ and samples from $R|_{R_{I}=1}$. [^9] Importantly, notice that $F$ does not get $I$ as input. We prove the following stronger lemma that finishes this part of the proof.
@@
We claim that Equation 8 follows by the assumption that $\mathsf{A}$ is a good distinguisher. Indeed, assume towards contradiction that this is not the case. Then, by a simple computation (and using the fact that $\mathsf{A}(R,Z_{R},W)\in\mathopen{}\mathclose{{}\left\{0,1}\right\}$) we get that
@@
and thus, using the triangular inequality we can write,
@@
which is in contradiction to our assumption on $\mathsf{A}$.

### v2

@@ 2 Proof Overview
We define the following random variables with respect to a random execution of 2.0.0.2. Let $X,Y$ be inputs of $\mathsf{A}$ and $\mathsf{B}$ to $\Pi$, respectively, and let $V_{A}$ and $V_{B}$ be their view in the interaction according to $\Pi$. Let $R,I_{1},\dots,I_{k},\widetilde{Y}$ and $\widehat{Y}$ be the random variables taking the values of $r,i_{1},\dots,i_{k},\widetilde{y}$ and $\widehat{y}$, respectively. Finally, let $\widehat{V}_{A}=(V_{A},R,\widehat{Y}_{-R})$ and $\widehat{V}_{B}=(V_{B},R,X_{R},I_{1},\dots,I_{k},\widetilde{Y},\widehat{Y})$ be the parties’ views in 2.0.0.2, and $O_{A},O_{B}$ their outputs, respectively.
It is evident by the definition of the protocol that erasure indeed occurs with probability $1/2$. Moreover, similarly to the correctness of the key-agreement protocol, and by the accuracy of $\Pi$, when there is no erasure, $O_{A}$ and $O_{B}$ are at most $\ell$ apart with probability at least $0.999$. Thus, it remains to prove that $\mathsf{B}$ cannot approximate $O_{A}$ when erasure occurs, and that $\mathsf{A}$ cannot learn whether it happened. To simplify the analysis in this proof overview, in the following, we assume that $\Pi$ is a $(\varepsilon,0)$-CDP protocol, [^5] While $(\varepsilon,0)$-CDP (that is, setting $\delta=0$) doesn’t make sense in the computational setting, for the sake of simplicity, we ignore this subtly in this informal overview. for a small constant $\varepsilon>0$.
Note that while $\mathsf{B}$ learns the same information on $\mathsf{A}$’s input in the protocol in both the erasure and non-erasure cases, $\mathsf{A}$’s output $O_{A}$ is different in each case; when there is no erasure, $O_{A}$ is predictable (up to an additive error $\ell$) given $\mathsf{B}$’s view $V_{B}$. When there is an erasure, we claim that the privacy guarantee of $\Pi$ implies that $O_{A}$, given $V_{B}$, is unpredictable up to an additive error $k^{1/2}\gg\ell$ (more details in Section 2.1). The “magic” of the protocol is in the fact that, for $\ell\ll n^{1/6}$, $\mathsf{B}$ can control $\mathsf{A}$’s output, and manipulate it to be unpredictable (from $\mathsf{B}$’s point of view), without $\mathsf{A}$ noticing the difference. Proving the latter property (hidden erasure) is the main technical challenge of this work, and we provide an overview of the proof in Section 2.2. By combining the two requirements on $k$, we obtain that our protocol is indeed ${\sf AWEC}$.
@@
Before discussing the security proof of our protocol, we would like to highlight an essential difference between the analysis of our protocol and [HaitnerMST22]’s ${\sf KA}$ protocol. While the primitive we construct here, ${\sf OT}$, is a stronger primitive than ${\sf KA}$, certain aspects of the proof become more manageable in our analysis. This is because in ${\sf OT}$, we need to consider an internal attacker, while in ${\sf KA}$ the security needs to hold against an external observer. In more detail, [HaitnerMST22] requires showing that security holds against an adversary who sees $R,X_{R}$, and $Y_{-R}$. This form of leakage creates a dependency between the seed $R$ and both the leakage on $X$, $X_{R}$, and the leakage of $Y$, $Y_{-R}$, which makes the analysis challenging. In contrast, in our setting, the attacker sees one of the vectors, $X$ or $Y$, entirely. For example, $\mathsf{A}$ gets to see $X,R$ and $Y_{-R}$, and therefore there is no dependency between the leakage on $X$ and $R$. This fact helps us to bypass this dependency issue and makes this part of our analysis more manageable.
2.1 Erasures do Erase
To see why “erasures” in 2.0.0.2 do erase the value of $O_{A}$ from $\mathsf{B}$’s point of view, let’s first consider a different version of 2.0.0.2, where in the erasure case, party $\mathsf{B}$, instead of changing $k$ random bits in $Y_{-R}$, changes all the bits to random ones. (Namely, in the erasure case, $\widehat{Y}_{-R}$ is a uniformly random string, independent of $X$). Now, the question is, can $\mathsf{B}$ now predict $O_{A}=\langle X_{-R},\widehat{Y}_{-R}\rangle$? From $\mathsf{B}$’s point of view, the privacy guarantee of the protocol $\Pi$ implies that $X_{-R}$ must be “unpredictable”, and in [McGregorMPRTV10, HaitnerMST22], it is called “strong Santha Vazirani (SV)” source. [McGregorMPRTV10] showed that the inner product of an $n$-size strong SV source with a random string, even if the string is known, is unpredictable up to an additive error of $\approx\sqrt{n}$, and [HaitnerMST22] extended this result to the computational case. Therefore, if $\mathsf{B}$ decides to erase in this version, it cannot estimate $O_{A}=\langle X_{-R},\widehat{Y}_{-R}\rangle$ with an error smaller than $\approx\sqrt{n}$, which is much larger than $\ell$, its estimation of $\langle X_{-R},\widehat{Y}_{-R}\rangle$ in the non-erasure case. The problem with this version is that $\mathsf{A}$ can distinguish between the erasure and non-erasure cases. Hence, in the actual protocol, the number of bits that $\mathsf{B}$ changes in the erasure case is set to some smaller parameter $k$. The point is that the same argument we applied to the previous version still holds; if $\mathsf{B}$ changes $k$ random values in $Y_{-R}$ (call this $k$-size set of indices by ${\cal H}$), then $O_{A}$ would be a linear function of $\langle X_{{\cal H}},\widehat{Y}_{{\cal H}}\rangle$, and since $\mathsf{B}$ cannot estimated $\langle X_{{\cal H}},\widehat{Y}_{{\cal H}}\rangle$ with error smaller than $\approx\sqrt{k}$ (from the same arguments above), it cannot do it for $O_{A}$ as well.
In more detail, assume towards a contradiction that there exists an efficient $\widehat{\mathsf{B}}$, that given $\widehat{V}_{B}$ and conditioned on $O_{B}=\bot$ (equivalently, on $b=1$) can approximate $O_{A}$ within a small distance. Namely, assume that
@@ 2.1 Erasures do Erase
To see why, let ¯=$\bar{{\cal H}}=\mathopen{}\mathclose{{}\left\{i\colon R_{i}=0}\right\}\setminus{\cal H}$, and let $T=\widehat{\mathsf{B}}(\widehat{V}_{B})-\langle X_{\bar{{\cal H}}},\widetilde{Y}_{\bar{{\cal H}}}\rangle$. We claim that $T$ is a good approximation of $\langle X_{{\cal H}},S\rangle$. Indeed,
$T-\langle X_{{\cal H}},S\rangle$ $=T-(\langle X_{-R},\widetilde{Y}_{-R}\rangle-\langle X_{\bar{{\cal H}}},\widetilde{Y}_{\bar{{\cal H}}}\rangle)$
$=(T+\langle X_{\bar{{\cal H}}},\widetilde{Y}_{\bar{{\cal H}}}\rangle)-\langle X_{-R},\widetilde{Y}_{-R}\rangle$
@@
Thus, we only need to show an algorithm $\widetilde{\mathsf{B}}$ that, given $V_{B},{\cal H},S,X_{-{\cal H}}$ as input, evaluates the value of $T$. The main issue left is that while $\widetilde{\mathsf{B}}$ only gets $V_{B}$, and to evaluate $T$ we need to execute $\widehat{B}$ that needs to get an entire view $\widehat{V}_{B}$ of $\mathsf{B}$ in 2.0.0.2. However, this can be easily simulated. Indeed, $\widetilde{\mathsf{B}}$ can simply sample $R,I_{1},\dots,I_{k}$ and $\widetilde{Y}$ condition on the values of ${\cal H},S$ (and $O_{B}=\bot$), to get the exact distribution of $\widehat{V}_{B}$.
2.2 Erasures are Hidden
This part of the proof is more involved and is our main new technical part. The goal is to show that $\widehat{\mathsf{A}}$ cannot tell when an erasure happens; equivalently, that $\widehat{\mathsf{A}}$ cannot distinguish $Y_{-R}$ from the noisy version $\widetilde{Y}_{-R}$ with too good probability. For this end, we show that an algorithm $\widehat{\mathsf{A}}$ that (given $V_{A}$) can distinguish $Y_{-R}$ from the noisy version $\widetilde{Y}_{-R}$ with a small but constant probability, can be used to break ${\sf CDP}$.
[…]
The first step is to notice that, by a rather standard hybrid argument, it is enough to show that $\widehat{\mathsf{A}}$ cannot distinguish with advantage $\Theta(1/k)$ between $Y$ and a vector $Y^{\prime}$ which is identical in all but one random coordinate. That is, let $I\leftarrow[n]$ be a random index, and let $Y^{(I)}=(Y_{1},\dots,Y_{I-1},-Y_{I},Y_{I+1},...,Y_{n})$ be the vector that is identical to $Y$ in all entries except the $I$-th one. Then it is enough to show that if
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A},R,Y_{-R})=1}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A},R,Y^{(I)}_{-R})=1}\right]\geq 1/k,$
@@ 2.2 Erasures are Hidden
We highlight that the key difficulty in proving the above lies in the fact that the distinguishing advantages are very small. Indeed, an algorithm $\widehat{A}$ that distinguishes $Y_{-R}$ from $Y^{(I)}_{-R}$ with a large enough constant advantage (proportional to $\varepsilon$) already contradicts the privacy requirement of $\Pi$. However, in our setup, the distinguishing advantage is much smaller than $\varepsilon$, which makes the task of contradicting $\varepsilon$-CDP challenging. We next emphasize two crucial points that make $\widehat{\mathsf{A}}$ useful (and our design for our AWEC protocol, 2.0.0.2, was guided by these two observations).
The importance of a random index
We note that it is crucial that the index $I$ is chosen at random and that it is unknown to $\widehat{\mathsf{A}}$. To see that, let $\mathsf{M}$ be a DP mechanism and let $Y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ be an input. We claim that it can be the case that an adversary can distinguish between $Y$ and $Y^{(i)}$ given $\mathsf{M}(Y)$ with advantage $\approx\varepsilon$, for any fixed index $i$. Indeed, it is possible that $\mathsf{M}(Y)$ outputs a noisy estimation of $Y_{i}$, which is correct with probability $\approx 1/2+\varepsilon$. Such an estimation will not contradict differential privacy, but will be useful to distinguish between $Y$ and $Y^{(i)}$. We resolve this issue by adding the noise in a random position. To prove that the existence of such noise (in a random position) is indeed hidden, we developed and proved Section 2.2 which is our main technical lemma.
The importance of a random subset
Next, we claim that it is crucial that $\widehat{\mathsf{A}}$ only gets $Y_{-R}$, a random subset of $Y$. Indeed, as in the previous example, the mechanism $\mathsf{M}(Y)$ can output a noisy estimation of the parity of the number of ones in $Y$. As in the previous example, such an estimation does not contradict DP, but is enough to distinguish $Y$ and $Y^{(I)}$ by considering the number of ones in the input compared to the estimation. We deal with such a global information by revealing to $\widehat{\mathsf{A}}$ only a subset of $Y$. This makes the global information not useful.
Thus, in the proof, we need to leverage the fact that $\widehat{\mathsf{A}}$ distinguishes $Y_{-R}$ from $Y^{(I)}_{-R}$ for a random index $I$ and a random subset $R$. We use such an algorithm $\widehat{\mathsf{A}}$ to construct an algorithm $\mathsf{G}$ that given $Y_{-I}=(Y_{1},\ldots,Y_{I-1},Y_{I+1},\ldots,Y_{n})$, predicts the value of $Y_{I}$. In more detail, we prove the following lemma, which is the technical core of our proof.
@@
Let $(Z,W)$ be jointly distributed random variables, let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $I\leftarrow[n]$. Let $\mathsf{A}$ be an efficient algorithm that satisfies
@@
Back to our setting, by taking $\mathsf{A}(R,Z_{R},W)=\widehat{\mathsf{A}}((1,\ldots,1)-R,Z_{R},W)$, $Z=Y$ and $W=V_{A}$ in Section 2.2, we get an algorithm $\mathsf{G}$ such that
@@
We highlight that the gap between the probability of $\mathsf{G}$ to output $Y_{I}$ to its probability to output $-Y_{I}$ is smaller than $1/k$, and that it is not enough to break DP alone. The point is that we have a bound on the probability of $\mathsf{G}$ to output $-Y_{I}$, and that in DP, we measure the difference in probabilities using multiplicative distance. Thus, when
@@
which is the reason we need to take $k\ll n^{1/3}$. In the rest of this section, we give a high-level proof sketch of Section 2.2.
2.3 Proof Overview for Section 2.2
In this part, we explain the ideas behind the proof of Section 2.2. We start by describing the algorithm $\mathsf{G}$.
@@
The construction of $\mathsf{G}$ leverages the fact that $\mathsf{A}$ succeed on a random index $I$ while only seeing a subset of $Y$. Recall that on input $(I,Z_{-I},W)$, the goal of $\mathsf{G}$ is to predict the value of $Z_{I}$. Towards this end, let $Z^{b}=(Z_{1},\dots,Z_{I-1},b,Z_{I+1},\dots,Z_{n})$ (hence, $Z\in\mathopen{}\mathclose{{}\left\{Z^{-1},Z^{1}}\right\}$). Algorithm $\mathsf{G}$ computes the following values: [^8] In the proof itself, $\mathsf{G}$ only estimates these values by sampling $R$. For simplicity, here we assume that $\mathsf{G}$ can compute them exactly.
@@
Importantly, we note that when $R_{I}=0$, $Z_{R}$ does not contain $Z_{I}$, and thus $\mathsf{G}$ can compute $\mu^{*}$, given its input. Finally, $\mathsf{G}$ outputs $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ if the following two conditions holds:
@@
Therefore, assuming that Item 1 and Item 2 hold, this conclude the proof of Section 2.2. We next briefly explain why Item 1 and Item 2 hold with the claimed probability.
@@
In the following, we will prove a stronger claim. That is, that for every fixed $Z$ and $W$, it holds that
@@
Finally, Equation 5 follows a more general well-known fact on the uniform distribution over random strings: For a random string $R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and a random index $I\leftarrow[n]$, the statistical distance between $R|_{R_{I}=0}$ and $R|_{R_{I}=1}$ is at most $1/\sqrt{n}$. This fact implies that the function $F$ cannot distinguish too well between samples from $R|_{R_{I}=0}$ and samples from $R|_{R_{I}=1}$. [^9] Importantly, notice that $F$ does not get $I$ as input. We prove the following stronger lemma that finishes this part of the proof.
@@
We claim that Equation 8 follows by the assumption that $\mathsf{A}$ is a good distinguisher. Indeed, assume, for the sake of contradiction, that this is not the case. Then, by a simple computation (and using the fact that $\mathsf{A}(R,Z_{R},W)\in\mathopen{}\mathclose{{}\left\{0,1}\right\}$) we get that
@@
and thus, using the triangular inequality, we can write,
@@
which contradicts our assumption on $\mathsf{A}$.
3 Related Works